# A note on generalized tensor CUR approximation for tensor pairs and tensor triplets based on the tubal product

### v3

@@ Abstract
In this note, we briefly present a generalized tensor CUR (GTCUR) approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ based on the tubal product (t-product). We use the tensor Discrete Empirical Interpolation Method (TDEIM) to do these extensions. We show how the TDEIM can be utilized to generalize the classical tensor CUR (TCUR) approximation, which acts only on a single tensor, to jointly compute the TCUR of two and three tensors. This approach can be used to sample relevant lateral/horizontal slices of one data tensor relative to one or two other data tensors. For some special cases, the Generalized TCUR (GTCUR) method is reduced to the classical TCUR for both tensor pairs and tensor triplets in a similar fashion as shown for the matrices.
@@ 1 Introduction
The matrix CUR (MCUR) approximation is a well-known technique for fast low-rank approximation of matrices [1]. It offers an interpretable approximation and a more compact data representation as it uses the actual columns/rows of the underlying data matrix to build the factor matrices. The generalized MCUR (GMCUR) approximation for matrix pairs and matrix triplets were presented in [2] and [3], respectively. Similar to the SVD, the MCUR is applied to a single matrix while the GMCUR considers a pair and a triplet of matrices and provides low-rank approximations of two/three matrices simultaneously. The key idea behind this generalization is using the same column indices for two given data matrices. To be more precise, let two data matrices ${\bf X}\in\mathbb{R}^{I_{1}\times I_{2}},\,(I_{1}\geq I_{2}),\,{\bf Y}\in\mathbb{R}^{I_{3}\times I_{2}},\,(I_{3}\geq I_{2})$ that have the same number of columns be given. The generalized SVD (GSVD) [4, 5], guarantees the following decompositions
@@
where $\alpha_{i}^{2}+\beta_{i}^{2}=1$ with the ratios $\beta_{i}/\alpha_{i}$ of nondecreasing order for $i=1,2,\ldots,I_{2}$. Here, ${\bf U}\in\mathbb{R}^{I_{1}\times I_{1}},\,{\bf V}\in\mathbb{R}^{I_{3}\times I_{3}}$ are orthogonal, while ${\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}}$ is a nonsingular matrix. As can be seen, the GSVD shares the common factor matrix ${\bf Z}$ between the given data matrices ${\bf X}$ and ${\bf Y}$. On the other hand, from the theory of the MCUR approximation, we know that by applying the Discrete Empirical Interpolation Method (DEIM) [6] to the basis ${\bf Z}$ (to be discussed in Section 3), we can sample appropriate column indices from the data matrices ${\bf X}$ and ${\bf Y}$. Since the same basis matrix ${\bf Z}$ is used within the DEIM, the same indices will be produced for sampling the columns of the matrix ${\bf X}$ and the matrix ${\bf Y}$. Besides, the matrices ${\bf U}$ and ${\bf V}$ can be used to sample row indices of the matrices ${\bf X}$ and ${\bf Y}$, respectively. This means that the selected row indices for these matrices can be different. As a result, for two matrices ${\bf X}$ and ${\bf Y}$ with the same number of columns, we can use the identical column indices for their MCUR approximations. The authors in [2] show that this idea can be potentially used in applications where one is interested in extracting the most discriminative features from one data set relative to another data set. It is also motivated for recovering data matrices perturbed with colored noise and subgroup discovery application. Following this idea, in [3], the authors consider triplet tensors $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ and use the restricted SVD [7] to build their common factor matrices. Then, the DEIM is applied to the common factor matrices to sample columns and rows. It is shown that this approach can be used to extract the most discriminative features from one data set relative to another data set. The GMCUR for tensor triplets is also motivated by the canonical correlation analysis and was successfully used for recovering datasets perturbed with colored noise, See [2, 2] for details.
@@
In this note, due to the applications of tensors [8, 9, 10], we focus on the tensor SVD (t-SVD) [11] that is defined based on the tubal product (t-product). The t-SVD has similar properties as the classical SVD. In particular, in contrast to the Tucker decomposition [12, 13] or Canonical polyadic decomposition [14, 15], its truncation provides the best low tubal rank approximation in the least-squares sense. The GSVD was generalized to tensors based on the t-SVD in [16, 17]. Motivated by the works [2, 3], we extend the tensor CUR (TCUR) approximation [18] to a pair and a triplet of tensors based on the t-product and we refer to as the generalized TCUR (GTCUR). As real-world datasets often have multidimensional structure, it is more favorable to generalize the ideas proposed in [2, 19] to tensors so that the structure of the datatensors are preserved. Indeed, we aim at achieving this target and present several generalizations of GMCUR to tensors. To this end, we use the tensor DEIM (TDEIM) index selection algorithm [20], which we recently developed to sample important lateral and horizontal slices, and this plays a key role in the development of the GTCUR method. Similar to the matrix case, the TDEIM helps to compute the same lateral/horizontal slices for a given pair or triplet of third order tensors. We remark that alternative sampling procedures, such as top tubal leverage scores [21, 18] or its sampling version, may be employed in place of the TDEIM, but because the TDEIM has been shown to give the optimal sample method [6], we adopt it in our research.
@@
The remainder of this note is organized as follows. We first present some basic tensor concepts in Section 2. The t-SVD is introduced in Section 2.1 with some intuitions behind this model. The matrix cross approximation is outlined in Section 3 and the Tensor Discrete Empirical Interpolation Method (TDEIM) as an extension of the DEIM method is presented in Section 4. In Sections 5 and 6, we show how the matrix cross approximation for matrix pairs/triplets can be generalized to tensor pairs and tensor triplets, respectively. The conclusion is given in Section 7.
@@ 2 Preliminaries
Before presenting the main materials, we first introduce the basic notations and definitions. An underlined bold capital case letter, a bold capital case letter, and a bold lower case letter are used to represent a tensor, a matrix, and a vector, respectively. For a third-order tensor $\underline{{\bf X}},$ we call the slices $\underline{{\bf X}}(:,:,k),\,\underline{{\bf X}}(:,j,:),\,\underline{{\bf X}}(i,:,:)$ frontal, lateral, and horizontal slices. For simplicity, sometimes for a frontal slice $\underline{\bf X}(:,:,i)$, we use the notation $\underline{\bf X}_{i}$. For a third-order tensor $\underline{\bf X},$ the fiber $\underline{{\bf X}}(i,j,:)$ is called a tube. The matrix’s component-wise complex conjugate is denoted by the symbol “conj”. To indicate a subset of a matrix or tensor, we use the MATLAB notations. For instance, for a given data matrix ${\bf X}$, by ${\bf X}(:,\mathcal{J})$ and ${\bf X}(\mathcal{I},:)$ we mean two matrices sampling a part of the rows and columns of the matrix ${\bf X}$, respectively, where $\mathcal{I}\subset\{1,2,\ldots,I_{2}\}$ and $\mathcal{J}\subset\{1,2,\ldots,I_{1}\}$. We need to present Definitions 1-5 to introduce the tensor SVD (t-SVD) model.
@@
(Transpose) The transpose of a tensor $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is denoted by $\underline{\mathbf{X}}^{T}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$ produced by applying the transpose to all frontal slices of the tensor $\underline{\mathbf{X}}$ and reversing the order of the transposed frontal slices from the second until the last one.
@@
The MP pseudoinverse of a tensor can also be computed in the Fourier domain and this is shown in Algorithm 2.
@@ 3       $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)={\rm pinv}\,(\widehat{{\bf X}}(:,:,i))$;
2.1 Tensor decompositions based on the t-product and tubal leverage-scores
@@
$\underline{{\bf X}}\approx\underline{\bf U}*\underline{\bf S}*\underline{\bf V}^{T},$
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf S}\in\mathbb{R}^{R\times R\times I_{3}},$ and $\underline{\bf V}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$. The tensors $\underline{\bf U}$ and $\underline{\bf V}$ are orthogonal, while the tensor $\underline{\bf S}$ is f-diagonal. The procedure of the computation of the t-SVD is presented in Algorithm 3.
@@ 1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
3       $[\widehat{\underline{\mathbf{U}}}\left({:,:,i}\right),\widehat{\underline{\bf S}}(:,:,i),\widehat{{\bf V}}(:,:,i)]={\rm Truncated\operatorname{-}svd}\,(\widehat{{\bf X}}(:,:,i),R)$;
@@
3 Matrix CUR (MCUR) approximation of a single matrix and matrix pairs
The Matrix CUR (MCUR), or cross approximation, is a well-known method for fast low-rank matrix approximation with interpretable factor matrices and linear computational complexity [1]. This technique proposes to choose the individual columns and rows of a matrix in order to preserve the original data matrix’s structure, such as negativity or sparsity. Let us describe it formally. Given ${\bf X}\in\mathbb{R}^{I_{1}\times I_{2}}$ be a give data matrix. The MCUR approximation seeks the approximation of the form ${\bf X}\approx{\bf C}{\bf U}{\bf R}$ where ${\bf C}={\bf X}(:,\mathcal{J}),\,{\bf R}={\bf X}(\mathcal{I},:)$ and $\mathcal{I}\subset\{1,2,\ldots,I_{1}\}$ and $\mathcal{J}\subset\{1,2,\ldots,I_{2}\}$. It is obvious that the best middle matrix is
@@ 3 Matrix CUR (MCUR) approximation of a single matrix and matrix pairs
A deterministic or random selection of the columns and rows is possible, with the option of obtaining additive or relative approximation errors. It is generally known that the columns or rows with the highest volume might provide virtually optimal solutions in a deterministic scenario [29]. The DEIM is another deterministic way to select a matrix’s columns or rows and relies on the top singular vectors [6]. The three probability distributions that are most typically used to choose columns and rows are uniform, length-squared, and leverage-score distributions. For additional details on these sampling techniques, see [21]. It has been shown that approximations with relative error accuracy that are more practical in practice can be produced by sampling columns with the leverage-score probability distribution [30].
The MCUR approximation of matrix pairs (matrices with the same number of columns) as the extension of the classical MCUR approximations were presented in [2]. The key idea of this generalization is to select columns from two matrices that share the same column indices in order to extract the features of one dataset that are most pertinent to the other. Here, the common factor matrix produced by the GSVD of the two matrices is used to sample the columns and rows using the Discrete Empirical Interpolation Method (DEIM) technique. The MCUR approximation of triplet matrices was also presented in [3] , with the goal of identifying pertinent characteristics of one data matrix in relation to two other matrices. The authors in this instance apply the RSVD to produce two common factor matrices, to which the DEIM is then applied to sample the column and row indices. We first discuss the DEIM extension to tensors based on the t-product in the next section before using it to extend the TCUR technique to tensor pairs and tensor triplets in Sections 5 and 6, respectively.
@@ 4 Tensor Discrete Empirical Interpolation Method (TDEIM) for lateral/horizontal slice sampling
We have recently generalized the DEIM method to the tensor case based on the t-product [20] that we call tensor DEIM (TDEIM). It has been experimentally shown that the TDEIM achieved the best sampling accuracy compared to the existing sampling techniques, such as top tubal leverage score sampling and uniform sampling. The TDEIM to select horizontal slices is presented in Algorithm 4. A key concept in the TDEIM is the interpolatory projector, as we now define it. For a given set of indices ${\bf s}\in\mathbb{R}^{R}$, the full tubal-rank tensor [^1] A tensor with linear independent lateral slices, for example, the tensor $\underline{\bf U}$ obtained from the t-SVD can be used. $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ and considering $\underline{\bf S}=\underline{\bf I}(:,{\bf s},:)\in\mathbb{R}^{I_{1}\times R\times I_{3}}$, as the selection tensor where $\underline{\bf I}\in\mathbb{R}^{R\times R\times I_{3}}$ is an identity tensor, we build the tensorial oblique projection operator defined as follows
@@
This means that the projection operator $\underline{\mathcal{P}}$ preserves the horizontal slices of $\underline{\bf G}$ specified by the index set ${\bf s}$. This justifies the name of interpolation. The TDEIM starts with selecting an index with the maximum Euclidean norm of the tubes in the first lateral slice of the basis, i.e. $\underline{\bf U}(:,1,:),$ to be the index of the first sampled horizontal slice. The subsequent indices are selected according to the indices with the maximum Euclidean norm of the tubes of the residual lateral slice that is computed by removing the direction of the tensorial interpolatory projection in the previous basis vectors from the subsequent one. To be more precise, let the indices ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\}$ have been already selected, and we want to select $s_{j}$. To do so, we compute the residual slice
@@
that will be used in the next theorem. Here, $\underline{\bf V}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ is a bais tensor for the subspace of horizontal slices of the original data tensor, $\underline{\bf Q}=\underline{\bf I}(:,{\bf s}_{j-1},:)$ is a tensor of some sampled lateral slices of the identity tensor specified by the index set ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\},\,\widehat{\underline{\bf Q}}={\rm fft}(\underline{\bf Q},[],3),\,\widehat{\underline{\bf V}}={\rm fft}(\underline{\bf V},[],3),\,\widehat{\underline{\bf Q}}_{i}=\widehat{\underline{\bf Q}}(:,:,i),$ and $\widehat{\underline{\bf V}}_{i}=\widehat{\underline{\bf V}}(:,:,i)$.
@@
[20] Suppose $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $1\leq R<\min(I_{1},I_{2})$. Assume that horizontal slice and lateral slice indices ${\bf p}$ and ${\bf q}$ give full tubal rank tensors $\underline{\bf C}=\underline{\bf X}(:,{\bf q},:)=\underline{\bf X}*\underline{\mathcal{Q}}$ and $\underline{\bf R}=\underline{\bf X}({\bf p},:,:)={\underline{\mathcal{P}}}*\underline{\bf X}$, where $\underline{\mathcal{P}}$ and $\underline{\mathcal{Q}}$ are tensorial interpolation projectors for horizontal and lateral slice sampling [^2] The tensorial interpolation projector for the lateral slice sampling is defined as $\underline{\mathcal{Q}}=\underline{{\bf V}}*(\underline{{\bf Q}}^{T}*\underline{{\bf V}})^{-1}*\underline{{\bf Q}}^{T}$, where $\underline{{\bf V}}$ is a basis tensor and $\underline{{\bf Q}}$ is the collection tensor., respectively with corresponding finite error constants $\tilde{\eta}_{p},\,\tilde{\eta}_{q}$ defined in (8) and set $\underline{\bf U}=\underline{{\bf C}}^{{\dagger}}*\underline{\bf X}*\underline{\bf R}^{{\dagger}}$. Then

### v4

@@ Abstract
In this note, we briefly present a generalized tensor CUR (GTCUR) approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ based on the tubal product (t-product). We use the tensor Discrete Empirical Interpolation Method (TDEIM) to do these extensions. We demonstrate how the TDEIM can be applied to extend the traditional tensor CUR (TCUR) approximation, which operates on a single tensor, to simultaneously compute the TCUR approximations for two or three tensors. This method allows for the sampling of relevant lateral or horizontal slices from one data tensor in relation to one or two other data tensors. In certain special cases, the Generalized TCUR (GTCUR) method simplifies to the classical TCUR approximations for both tensor pairs and tensor triplets, akin to the process shown for matrices.
@@
[…]
[label3]organization=Razi University, Kermanshah, Iran,
@@ 1 Introduction
The matrix CUR (MCUR) approximation is a well-known technique for fast low-rank approximation of matrices [17]. It offers an interpretable approximation and a compact data representation as it uses the actual columns/rows of the underlying data matrix to build the factor matrices. The generalized MCUR (GMCUR) approximation for matrix pairs and matrix triplets were presented in [13] and [15], respectively. Like the SVD, the MCUR is utilized for a single matrix, whereas the GMCUR addresses a pair or triplet of matrices, offering low-rank approximations for two or three matrices at the same time. The central concept behind these generalizations is to employ the same column or row indices for the two data matrices in their low-rank approximations. To be more precise, let two data matrices ${\bf X}\in\mathbb{R}^{I_{1}\times I_{2}},\,(I_{1}\geq I_{2}),\,{\bf Y}\in\mathbb{R}^{I_{3}\times I_{2}},\,(I_{3}\geq I_{2})$ that have the same number of columns be given. The generalized SVD (GSVD) [37, 31], guarantees the following decompositions
@@
where $\alpha_{i}^{2}+\beta_{i}^{2}=1$ with the ratios $\beta_{i}/\alpha_{i}$ of nondecreasing order for $i=1,2,\ldots,I_{2}$. Here, ${\bf U}\in\mathbb{R}^{I_{1}\times I_{1}},\,{\bf V}\in\mathbb{R}^{I_{3}\times I_{3}}$ are orthogonal, while ${\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}}$ is a nonsingular matrix. As can be seen, the GSVD shares the common factor matrix $\bf Z$} between the given data matrices $\bf X$} and $\bf Y$}. On the other hand, from the theory of the MCUR approximation, we know that by applying the Discrete Empirical Interpolation Method (DEIM) [33] to the basis $\bf Z$} (to be discussed in Section 3), we can sample appropriate column indices from the data matrices $\bf X$} and $\bf Y$}. Since the same basis matrix $\bf Z$} is used within the DEIM, the same indices will be produced for sampling the columns of the matrix $\bf X$} and the matrix $\bf Y$}. Besides, the matrices $\bf U$} and $\bf V$} can be used to sample row indices of the matrices $\bf X$} and $\bf Y$}, respectively. This means that the selected row indices for these matrices can be different. As a result, for two matrices $\bf X$} and $\bf Y$} with the same number of columns, we can use the identical column indices for their MCUR approximations. In [13] , the authors demonstrate that this concept can be effectively applied in scenarios where the goal is to identify the most distinguishing features from one dataset in relation to another. It is also relevant for recovering data matrices affected by colored noise and for subgroup discovery applications. Building on this idea, the authors in [15] examine triplet matrices $({\bf X},{\bf Y},{\bf Z})$ and employ restricted SVD [38] to construct their common factor matrices. They then apply the DEIM to these common factor matrices to sample both columns and rows. This method has been shown to effectively extract the most discriminative features from one dataset compared to two others. The GMCUR for matrix triplets is also inspired by canonical correlation analysis and has proven successful in recovering datasets influenced by colored noise; further details can be found in [13, 15].
@@
In this note, due to the applications of tensors [11, 6, 7], we focus on the tensor SVD (t-SVD) [24] that is defined based on the tubal product (t-product). The t-SVD has similar properties as the classical SVD. In particular, in contrast to the Tucker decomposition [35, 36] or Canonical Polyadic Decomposition [20, 21], its truncation provides the best low tubal rank approximation in the least-squares sense. The GSVD was generalized to tensors based on the t-SVD in [19, 40]. Inspired by the works in [13, 15], we extend the tensor CUR (TCUR) approximation [34] to accommodate pairs and triplets of tensors based on the t-product, which we refer to as the generalized TCUR (GTCUR) approximation. Given that real-world datasets frequently exhibit multidimensional structures, it is beneficial to adapt the concepts introduced in [13, 14] for tensors, ensuring that the integrity of the data tensor structures is maintained. Our objective is to achieve this goal, and we present several generalizations of GMCUR for tensors. To facilitate this, we utilize the tensor DEIM (TDEIM) index selection algorithm [2], which we recently developed to sample significant lateral and horizontal slices, playing a crucial role in the GTCUR method’s development. Like the matrix case, the TDEIM facilitates the computation of identical lateral/horizontal slices for a specified pair or triplet of third-order tensors. It is worth noting that alternative sampling methods, such as top tubal leverage scores [28, 34] or their sampling variants, could be used instead of the TDEIM. However, since the TDEIM has been demonstrated to provide the optimal sampling method [33, 2], we have chosen to adopt it in our research.
@@
The remainder of this note is organized as follows. We first present some basic tensor concepts in Section 2. The matrix cross approximation is outlined in Section 3 and the Tensor Discrete Empirical Interpolation Method (TDEIM) as an extension of the DEIM method is presented in Section 4. In Sections 5 and 6, we show how the matrix cross approximation for matrix pairs/triplets can be generalized to tensor pairs and tensor triplets, respectively. The numerical experiments are presented in Section 7. The conclusion is given in Section 8.
@@ 2 Preliminaries
Prior to introducing the main materials, we will first outline the fundamental notations and definitions. An underlined bold capital case letter, a bold capital case letter, and a bold lower case letter are used to represent a tensor, a matrix, and a vector, respectively. For a third-order tensor $\underline{{\bf X}},$ we call the slices $\underline{{\bf X}}(:,:,k),\,\underline{{\bf X}}(:,j,:),\,\underline{{\bf X}}(i,:,:)$ frontal, lateral, and horizontal slices. For simplicity, sometimes for a frontal slice $\underline{\bf X}(:,:,i)$, we use the notation $\underline{\bf X}_{i}$. For a third-order tensor $\underline{\bf X},$ the fiber $\underline{{\bf X}}(i,j,:)$ is called a tube. The matrix’s component-wise complex conjugate is denoted by the symbol “conj”. To indicate a subset of a matrix or tensor, we use the MATLAB notations. For instance, for a given data matrix $\bf X$}, by ${\bf X}(:,\mathcal{J})$ and ${\bf X}(\mathcal{I},:)$ we mean two matrices sampling a part of the rows and columns of the matrix $\bf X$}, respectively, where $\mathcal{I}\subset\{1,2,\ldots,I_{2}\}$ and $\mathcal{J}\subset\{1,2,\ldots,I_{1}\}$. For example, let
$A=\begin{bmatrix}1&2&3&4\\ 5&6&7&8\\ 9&10&11&12\\ 13&14&15&16\\ 17&18&19&20\end{bmatrix}\in\mathbb{R}^{5\times 4}$
and $\mathcal{I}=\{1,2\}\subset\{1,2,3,4,5\},\,\mathcal{J}=\{2,4\}\subset\{1,2,3,4\}$, then
$A(\mathcal{I},\mathcal{J})=\begin{bmatrix}2&4\\ 6&8\end{bmatrix}$.
We will employ the same notation when sampling indices of tensor modes.
To introduce the tensor SVD (t-SVD) model, we must first present Definitions 1-5.
@@
(Transpose) Let $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be a given tensor. Then the transpose of the tensor $\underline{\mathbf{X}}$ is denoted by $\underline{\mathbf{X}}^{T}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$, which is constructed by transposing all its frontal slices and then reversing the order of transposed frontal slices $2$ through $I_{3}$.
@@
The MP pseudoinverse of a tensor can also be computed in the Fourier domain and this is shown in Algorithm 2. The inverse of a tensor $\underline{\bf X}$, denoted by $\underline{\bf X}^{-1}$ is a special case of MP for which we have $\underline{\bf X}^{-1}*\underline{\bf X}=\underline{\bf X}*\underline{\bf X}^{-1}=\underline{\bf I}$.
@@ 3 $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)={\rm pinv}\,(\widehat{{\bf X}}(:,:,i))$;
$\underline{{\bf X}}=\underline{\bf U}*\underline{\bf S}*\underline{\bf V}^{T},$
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}},\,\underline{\bf S}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},$ and $\underline{\bf V}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$. The tensors $\underline{\bf U}$ and $\underline{\bf V}$ are orthogonal, while the tensor $\underline{\bf S}$ is f-diagonal. The procedure of the computation of the t-SVD is presented in Algorithm 3. A truncated T-SVD of tubal rank $R$ is obtained as follows:
$\underline{{\bf X}}\approx\underline{\bf U}_{R}*\underline{\bf S}_{R}*\underline{\bf V}_{R}^{T},$
where $\underline{\bf U}_{R}=\underline{\bf U}(:,1:R,:)\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf V}_{R}=\underline{\bf V}(:,1:R,:)\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ and $\underline{\bf S}_{R}=\underline{\bf S}(1:R,1:R,:)\in\mathbb{R}^{R\times R\times I_{3}}$.
@@ 1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
3 $\widehat{\underline{\mathbf{U}}}\left({:,:,i}\right),\widehat{\underline{\bf S}}(:,:,i),\widehat{{\bf V}}(:,:,i)]={\rm Truncated\operatorname{-}svd}\,(\widehat{{\bf X}}(:,:,i),R)$;
@@
3 Matrix CUR (MCUR) approximation of a single matrix, matrix pairs and matrix triples
A well-known technique for fast low-rank approximation is the Matrix CUR (MCUR) method, or cross approximation [17]. Its strength lies in selecting a small subset of the original matrix’s columns and rows to form an interpretable factorization. This approach maintains the matrix’s salient features, e.g. like sparsity and non-negativity, at linear computational cost. Let us describe it formally. For a given data matrix ${\bf X}\in\mathbb{R}^{I_{1}\times I_{2}}$ , the MCUR method seeks the approximation of the form ${\bf X}\approx{\bf C}{\bf U}{\bf R}$ where ${\bf C}={\bf X}(:,\mathcal{J}),\,{\bf R}={\bf X}(\mathcal{I},:)$ and $\mathcal{I}\subset\{1,2,\ldots,I_{1}\}$ and $\mathcal{J}\subset\{1,2,\ldots,I_{2}\}$. It is obvious that the best middle matrix is
@@ 3 Matrix CUR (MCUR) approximation of a single matrix, matrix pairs and matrix triples
It is possible to select columns and rows either deterministically or randomly, allowing for the option to achieve either additive or relative approximation errors. It is widely recognized that in a deterministic context, selecting columns or rows with the highest volume can lead to nearly optimal solutions [16]. The DEIM is another deterministic approach for selecting a matrix’s columns or rows, which is based on the leading singular vectors [33].. The three most commonly utilized probability distributions for column and row selection are uniform, length-squared, and leverage-score distributions. For further information on these sampling methods, refer to [28]. Research has demonstrated that using the leverage-score probability distribution for sampling columns can yield more practical approximations with relative error accuracy [10].
The MCUR approximation for matrix pairs (matrices that have the same number of columns) was introduced as an extension of classical MCUR approximations in [13]. The key idea of this generalization involves selecting columns from two matrices that share identical column indices to identify the features of one dataset that are most relevant to the other. In this approach, the common factor matrix obtained through the Generalized Singular Value Decomposition (GSVD) of the two matrices is utilized to sample the columns using the DEIM technique. Additionally, the MCUR approximation for triplet matrices was discussed in [15], aiming to uncover relevant characteristics of one data matrix in relation to two others. In this case, the authors employed the Randomized Singular Value Decomposition (RSVD) to generate two common factor matrices, after which DEIM is applied to sample both column and row indices. We will first explore the extension of DEIM to tensors based on the t-product in the next section, then we use it to extend the TCUR approximation to tensor pairs and tensor triplets in Sections 5 and 6, respectively.
@@ 4 Tensor Discrete Empirical Interpolation Method (TDEIM) for lateral/horizontal slice sampling
TCUR approximates a tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ as $\underline{\bf X}\approx\underline{\bf X}(:,\mathcal{I},:)*\underline{\bf U}*\underline{\bf X}(\mathcal{J},:,:)$, where $\underline{\bf X}(:,\mathcal{I},:)$ and $\underline{\bf X}(\mathcal{J},:,:)$ are actual lateral and horizontal slices of the original tensor and $\underline{\bf U}=\underline{\bf X}(:,\mathcal{I},:)^{\dagger}*\underline{\bf X}*\underline{\bf X}(\mathcal{J},:,:)^{\dagger}$, see Figure 1 for visual illustration and we refer to the review paper [1] for more details on different types of TCUR methods. Determining a suitable selection of lateral and horizontal slices is a complex task. Recently, we extended the DEIM method to tensors using the t-product [2], which we refer to as tensor DEIM (TDEIM). Experimental results demonstrate that TDEIM provides superior sampling accuracy compared to existing methods, including top tubal leverage score sampling and uniform sampling. The process for using TDEIM to select horizontal slices is outlined in Algorithm 4.
Appropriate section of lateral and horizontal slices is a challenging problem. We have recently generalized the DEIM method to the tensor case based on the t-product [2] that we call tensor DEIM (TDEIM). It has been experimentally shown that the TDEIM achieved the best sampling accuracy compared to the existing sampling techniques, such as top tubal leverage score sampling and uniform sampling. The process for using TDEIM to select horizontal slices is outlined in Algorithm 4.
Figure 1: Tensor CUR approximation based on slice sampling [34, 1].
A fundamental concept of the TDEIM is the interpolatory projector, which we will define now. For a given set of indices ${\bf s}\in\mathbb{R}^{R}$, the full tubal-rank tensor [^1] A tensor with linear independent lateral slices, for example, the tensor $\underline{\bf U}$ obtained from the t-SVD can be used. $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ and considering $\underline{\bf S}=\underline{\bf I}(:,{\bf s},:)\in\mathbb{R}^{I_{1}\times R\times I_{3}}$, as the index selection tensor [^2] It is called index selection tensor because $\underline{\bf X}({\bf s},:,:)=\underline{\bf S}^{T}*\underline{\bf X}.$. where $\underline{\bf I}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is an identity tensor, we build the tensorial oblique projection operator defined as follows
@@
This means that the projection operator $\underline{\mathcal{P}}$ preserves the horizontal slices of $\underline{\bf G}$ specified by the index set $\bf s$}. This justifies the name of interpolation. The TDEIM starts with selecting an index with the maximum Euclidean norm of the tubes in the first lateral slice of the basis, i.e. $\underline{\bf U}(:,1,:),$ and assign it as the index of the first sampled horizontal slice. The subsequent indices are selected according to the indices with the maximum Euclidean norm of the tubes of the residual lateral slice that is computed by removing the direction of the tensorial interpolatory projection in the previous basis vectors from the subsequent one. To be more precise, let the indices ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\}$ have been already selected, and we want to select $s_{j}$. To do so, we compute the residual slice
@@
that will be used in the next theorem. Here, $\underline{\bf V}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ is a basis tensor for the subspace of horizontal slices of the original data tensor, $\underline{\bf Q}=\underline{\bf I}(:,{\bf s}_{j-1},:)$ is a tensor of some sampled lateral slices of the identity tensor specified by the index set ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\},\,\widehat{\underline{\bf Q}}={\rm fft}(\underline{\bf Q},[],3),\,\widehat{\underline{\bf V}}={\rm fft}(\underline{\bf V},[],3),\,\widehat{\underline{\bf Q}}_{i}=\widehat{\underline{\bf Q}}(:,:,i),$ and $\widehat{\underline{\bf V}}_{i}=\widehat{\underline{\bf V}}(:,:,i)$.
@@
[2] Suppose $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $1\leq R<\min(I_{1},I_{2})$. Assume that horizontal slice and lateral slice indices $\bf p$} and $\bf q$} give full tubal rank tensors $\underline{\bf C}=\underline{\bf X}(:,{\bf q},:)=\underline{\bf X}*\underline{\mathcal{Q}}$ and $\underline{\bf R}=\underline{\bf X}({\bf p},:,:)={\underline{\mathcal{P}}}*\underline{\bf X}$, where $\underline{\mathcal{P}}$ and $\underline{\mathcal{Q}}$ are tensorial interpolation projectors for horizontal and lateral slice sampling [^3] The tensorial interpolation projector for the lateral slice sampling is defined as $\underline{\mathcal{Q}}=\underline{{\bf V}}*(\underline{{\bf Q}}^{T}*\underline{{\bf V}})^{-1}*\underline{{\bf Q}}^{T}$, where $\underline{{\bf V}}$ is a basis tensor and $\underline{{\bf Q}}$ is the index selection tensor. The same can be defined for horizontal slice selection where a basis tensor $\underline{\bf U}$ is used., respectively with corresponding finite error constants $\tilde{\eta}_{p},\,\tilde{\eta}_{q}$ defined in (8) and set $\underline{\bf U}=\underline{{\bf C}}^{{\dagger}}*\underline{\bf X}*\underline{\bf R}^{{\dagger}}$. Then